The short and long hands of a clock are 4 cm and 6 cm respectively. Find the sum of the distances travelled by their tips in two days.

Grade:12th Pass

3 Answers

Raju Garg
12 years ago

area of short hand = 22/7*4 =88/7

total area in two days = 88/7*4=362/7

area of long hand = 22/7*6=132/7

total area in two days=132/7*48=6336/7

total area covered in two days = 6336/7+362/7=6698/7 = 956.85

Swag buddy
6 years ago
Round travel by show than in 2 days is equal to 4 and round travelled by long hand in 2 days is equal to 48 there for distance travelled by tips of the hands in 2 days will be there circumference is that4×2π4+48×2π6=1911.71
Swag buddy
6 years ago
Round travel by short hand*( it was miss printed in the previous answer) in 2 days is equal to 4 and round travelled by long hand in 2 days is equal to 48 there for distance travelled by tips of the hands in 2 days will be there circumference is that4×2π4+48×2π6=1911.71